OpenStudy (anonymous):

Solve 7.1x – 2.1 = 6.6x – 1.8.

OpenStudy (anonymous):

7.1x - 6.6x = 2.1 - 1.8 0.5x = 0.3 x = 0.3/ 0.5 Therefore, x = 3/5